OpenStudy (anonymous):

Mathnificent is a math tutoring service. The first year it was open, it had 5 students. If 3 new students enroll each year, which of the following equations represents how many students will be enrolled in x years? y = 5x - 3 y = 3x + 5 y = 5x + 3 y = 3x - 5

Directrix (directrix):

At time zero, there were 5 students. So, taking the first equation, y = 5x -3, then at time zero, y = 5*0 - 3 = -3 which is NOT correct because at time zero, there are five students.

Directrix (directrix):

@katieekazoo Evaluate y = 3x + 5 for x = 0 and post what you get.